OpenStudy (anonymous):

surds: rationalise and simplify 1/square root of 2 divide by 1/square root of 2 - 1

OpenStudy (anonymous):

\[1\div \sqrt{2}\div1\div \sqrt{2}-1\]

OpenStudy (anonymous):

need working

OpenStudy (anonymous):

\[\frac{1}{\sqrt{2}} \div \frac{1}{\sqrt{2 - 1}} = \frac{\sqrt{2-1}}{\sqrt{2}} = \frac{1}{\sqrt{2}}\]

OpenStudy (anonymous):

so, you just divide two fractions (division = flip one and multiply) then 2-1 = 1, square root of 1 is 1

OpenStudy (anonymous):

or, just do sqrt(2-1) = 1 straight away